# Breaking the Dimensional Barrier: Dynamic Portfolio Choice with Parameter Uncertainty via Pontryagin Projection

### v2

@@ 3.1 PG–DPO as stochastic gradient ascent and a conditional BPTT–PMP correspondence
Theorem 2 (BPTT–PMP correspondence (conditional on $\theta$, including second-adjoint blocks; uniform on compacts)).
Fix $\theta\in\Theta$ and assume standard regularity conditions ensuring (i) well-posedness of the $\theta$-conditional forward SDE (45)–(46) under the $\theta$-blind policy $\pi_{\varphi}$, and (ii) well-posedness of the associated $\theta$-conditional stochastic maximum principle, including the wealth-component first adjoint and the relevant blocks of the second adjoint needed for the portfolio Hamiltonian gradient. Let $(p_{t}^{\theta},p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote these continuous-time Pontryagin objects under $\pi_{\varphi}$. In smooth Markov regimes, they coincide (along the controlled state process) with $(V_{x},V_{xx},V_{xy})$ for the $\theta$-conditional value function.
Let $(p_{k},p_{x,k},p_{y,k})$ be the discrete pathwise quantities computed by BPTT for the Euler discretization with step $\Delta t$, as defined in (51)–(52). Let $(p_{k}^{\Delta t,\theta},p_{x,k}^{\Delta t,\theta},p_{y,k}^{\Delta t,\theta})$ be their adapted projections onto $G_{t_{k}}^{\theta}$ as defined above, and let the associated discrete-time BSDE martingale coefficients be defined via the one-step conditional $L^{2}$ projection on each step. Then, as $\Delta t\to 0$, the projected (adapted) BPTT-induced discrete adjoints converge to their continuous-time counterparts in an appropriate mean-square sense (along trajectories). Moreover, for any compact set $K\subset\Theta$, the constants in the convergence bounds can be chosen uniformly for all $\theta\in K$.
Proof.
See Appendix C. ∎
@@
Across $\theta\sim q$, these Pontryagin objects form a $\theta$-indexed family. Baseline PG–DPO trains against the ex–ante objective (48) by repeatedly sampling $\theta\sim q$ inside the simulator, while the deployable policy remains $\theta$-blind.
@@
3.2 Projected PG–DPO under latent $\theta$: $q$-aggregated projection, diagnostics, and a residual-based policy-gap bound
@@
Stage 2 is a projection step. Given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ (from stage 1), we estimate $\theta$-conditional Pontryagin adjoint objects under frozen $\theta\sim q$ by BPTT/Monte Carlo, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated stationarity condition derived in Section 2.2. Equivalently, Stage 2 can be viewed as a plug-in approximation of one application of the population projection map $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ (defined below) evaluated at $\pi=\pi^{\mathrm{warm}}$ on the working domain. The key point is that the aggregated first-order condition is affine in the portfolio control; hence it induces a statewise linear system and, on a suitable working domain, a concrete projection map from (estimated) adjoint objects to a portfolio rule.
@@
Working domain and norms. Fix a measurable working state domain $D\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ (e.g. a training/evaluation band) and a reference measure $\mu$ on $D$ (e.g. an empirical state distribution induced by rollouts). For $h:D\to\mathbb{R}^{n}$ we write
$\|h\|_{L^{2}(\mu)}:=\Big(\int_{D}\|h(z)\|^{2}\,\mu(dz)\Big)^{1/2},\hskip 18.49988ptz=(t,x,y),$
@@
$\|f\|_{L^{2}(q\otimes\mu)}:=\bigg(\int_{\Theta}\int_{D}\|f^{\theta}(z)\|^{2}\,\mu(dz)\,q(d\theta)\bigg)^{1/2}.$ (53)
@@
By Theorem 1, any locally optimal interior deployable $\theta$-blind policy $\pi^{\star,\mathrm{blind}}$ for the fixed-$q$ ex–ante problem satisfies the $q$-aggregated stationarity condition (15). In the portfolio Hamiltonian (10), this stationarity is equivalent to a statewise linear system and hence to the projected form (20) on the working domain (under invertibility of the aggregated curvature term). P–PGDPO constructs a practical approximation of this projection by estimating the relevant aggregated Pontryagin objects under a fixed warm-up policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$.
Fix a query state $z=(t,x,y)\in D$ and a frozen parameter $\theta$. We simulate trajectories under $\pi^{\mathrm{warm}}$ and compute discrete-time adjoint blocks by autodiff/BPTT on the Euler simulator; averaging over $M_{\mathrm{MC}}$ trajectories yields Monte Carlo estimates
@@
Whenever $\widehat{A}_{t}(t,x,y)$ is invertible on $D$, we obtain the mixed-moment projected policy
@@
Stabilized inversion and projection diagnostics (implementation). In practice, finite-sample estimates can make $\widehat{A}_{t}$ ill-conditioned or even singular on parts of $D$. Our default implementation therefore uses standard stabilizers (Appendix E.5), including: (i) ridge regularization $\widehat{A}_{t,\lambda}:=\widehat{A}_{t}+\lambda I_{d}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\lambda}:=-\widehat{A}_{t,\lambda}^{-1}\widehat{G}_{t}^{\mathrm{mix}}$, (ii) condition-number or eigenvalue diagnostics to skip/clip projection on unreliable states (falling back to $\pi^{\mathrm{warm}}$), and (iii) residual-form projection (Section F.1) to reduce variance. The analysis below is stated for the population projection $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ and for small estimator perturbations, which is precisely the regime enforced by these stabilizers.
Remark (decoupled aggregation heuristic). The mixed-moment aggregation (57)–(58) is the object appearing in the exact $q$-aggregated first-order condition. In some high-dimensional experiments, we additionally consider a variance-reduction heuristic that replaces mixed moments by products of empirical means, e.g. $\mathbb{E}_{\theta}[\widehat{p}_{t}^{\theta}\,b(y,\theta)]\approx\mathbb{E}_{\theta}[\widehat{p}_{t}^{\theta}]\cdot\mathbb{E}_{\theta}[b(y,\theta)]$, and similarly for the other terms. All theoretical statements in this subsection are for the mixed-moment projection (59).
@@
To connect (59) to a locally optimal deployable $\theta$-blind policy, we measure how well the warm-up policy satisfies the population mixed-moment aggregated stationarity. Let $(A_{\pi},G_{\pi}^{\mathrm{mix}})$ denote the mixed-moment $q$-aggregated projection inputs induced by a policy $\pi$ (i.e. the objects in (18) evaluated using the $\theta$-conditional Pontryagin adjoint blocks generated by $\pi$). Define the warm-up aggregated stationarity residual on $D$ by
$r^{\mathrm{warm}}_{\mathrm{FOC,mix}}(t,x,y):=A_{\pi^{\mathrm{warm}}}(t,x,y)\,\pi^{\mathrm{warm}}(t,x,y)+G_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}(t,x,y),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\big\|r^{\mathrm{warm}}_{\mathrm{FOC,mix}}\big\|_{L^{2}(\mu)}.$ (60)
Note that $r^{\mathrm{warm}}_{\mathrm{FOC,mix}}=0$ is equivalent to $\pi^{\mathrm{warm}}$ being a fixed point of $T$ on $D$. More generally, the population projection admits the residual form
$T(\pi^{\mathrm{warm}})\;=\;\pi^{\mathrm{warm}}\;-\;A_{\pi^{\mathrm{warm}}}^{-1}\,r^{\mathrm{warm}}_{\mathrm{FOC,mix}}.$ (61)
In practice we monitor the empirical analogue $\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{G}_{t}^{\mathrm{mix}}$, and use it as a diagnostic to flag states where projection is unreliable (cf. Appendix E.5). (Equivalently, $\widehat{\pi}^{\mathrm{agg,mix}}=\pi^{\mathrm{warm}}-\widehat{A}_{t}^{-1}\widehat{r}^{\mathrm{warm}}_{\mathrm{FOC,mix}}$, or its stabilized variant, which motivates the residual-form implementation.)
@@
A technical point is that a global small-gain condition of the form $C_{1}<1$ can be overly restrictive. Following the slab-wise philosophy in our prior PGDPO analysis (e.g. Huh et al. (2025b, Appendix B)), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab. Concretely, assume $D$ carries a time coordinate and fix a partition $0=t_{0}<t_{1}<\cdots<t_{K}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$. Let
$D_{k}:=D\cap([t_{k-1},t_{k}]\times\mathcal{S}),\hskip 18.49988pt\mu_{k}:=\mu|_{D_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
We write $T(\pi):=-A_{\pi}^{-1}G_{\pi}^{\mathrm{mix}}$ for the (population) $q$-aggregated projection map. Theorem 3 below shows that, under a mild slab-wise local stability regime (i.e. a short-time contraction of $T$ on each $D_{k}$), small residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal deployable $\theta$-blind policy, up to discretization/Monte Carlo error. The proof combines a projection-map stability bound (Appendix D.1) with a slab-wise closure (Appendix D.2), in the same spirit as the slab analyses used in Huh et al. (2025b).
Stage 1: PG–DPO
@@
Stage 2: $q$-aggregated projection
@@
Figure 1: Two-stage pipeline of Section 3. A detailed schematic (including the coupling mechanisms and implementation notes) is provided in Appendix F, Fig. LABEL:fig:sec3-pipeline-sideways.
Theorem 3 (Residual-based ex–ante $\theta$-blind policy-gap bound for P–PGDPO (mixed-moment, deployable, slab-wise local)).
Assume the uniform invertibility/stability conditions of Proposition 1 (Appendix D.1) hold on $D$ for the relevant aggregated curvature terms and for the estimator perturbations constructed under $\pi^{\mathrm{warm}}$ (in particular, that $\widehat{A}_{t}$ remains a small perturbation of $A_{\pi^{\mathrm{warm}}}$ on $D$ under the adopted stabilizers).
Let $\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem. Assume there exists a neighborhood $\mathcal{U}$ of $\pi^{\star,\mathrm{blind}}$ in $L^{2}(\mu)$ such that for all $\pi\in\mathcal{U}$,
$\|A_{\pi}^{-1}\|_{L^{\infty}(D)}\leq\kappa,\hskip 18.49988pt\|G_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(D)}\leq M_{G},$
and assume the slab-wise Lipschitz gain of Appendix D.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{G}>0$ such that for every slab $D_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
@@
$\rho(\tau):=\big(\kappa\bar{L}_{G}+\kappa^{2}M_{G}\bar{L}_{A}\big)\tau^{1/2},\hskip 18.49988pt\rho_{*}:=\max_{1\leq k\leq K}\rho(\tau_{k}).$ (64)
Assume the slab partition is chosen so that $\rho_{*}<1$.
@@
$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\rho_{*}\kappa}{1-\rho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$ (65)

### v3

@@ 3.1 DPO via stochastic gradient ascent and a conditional BPTT–PMP correspondence
Theorem 2 (BPTT–BSDE/PMP correspondence (conditional on $\theta$, including second-adjoint blocks; uniform on compacts)).
Fix a compact $\Theta_{0}\subset\Theta$ and a $\theta$-blind policy network $\pi_{\varphi}$. Assume Assumption 1 holds for $\pi=\pi_{\varphi}$ and every $\theta\in\Theta_{0}$, and that the discrete simulator used for BPTT preserves positivity as in Remark 4.
For each $\theta\in\Theta_{0}$, let $(p_{t}^{\theta},z_{t}^{\theta},\tilde{z}_{t}^{\theta})$ denote the wealth-component adjoint BSDE solution (15) under $\pi_{\varphi}$, and let $(p_{x,t}^{\theta},p_{y,t}^{\theta})$ denote the blocks (16) (equivalently, the relevant blocks of the second adjoint along the controlled trajectory).
Let $(X_{k},Y_{k})_{k=0}^{N}$ be the Euler simulator on a grid $(t_{k})_{k=0}^{N}$ with step size $\Delta t$, and let $(p_{k},p_{x,k},p_{y,k})$ be the (pathwise, generally non-adapted) BPTT adjoint blocks (45)–(46). Define their adapted projections by conditional expectations under the simulator filtration $\mathcal{G}_{t_{k}}^{\theta}$:
$p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid\mathcal{G}_{t_{k}}^{\theta}].$
Let $(Z_{k}^{\theta},\tilde{Z}_{k}^{\theta})$ be the one-step conditional $L^{2}$ projection coefficients of $p_{k+1}$ onto $\mathrm{span}\{\Delta W_{k},\Delta W_{k}^{Y}\}$ given $\mathcal{G}_{t_{k}}^{\theta}$ (so that the orthogonal residual is $L^{2}$-uncorrelated with the increments). Form the piecewise-constant interpolations on $t_{k},t_{k+1})$:
$p_{t}^{\Delta t,\theta}:=p_{k}^{\Delta t,\theta},\qquad p_{x,t}^{\Delta t,\theta}:=p_{x,k}^{\Delta t,\theta},\qquad p_{y,t}^{\Delta t,\theta}:=p_{y,k}^{\Delta t,\theta},\qquad z_{t}^{\Delta t,\theta}:=Z_{k}^{\theta},\qquad\tilde{z}_{t}^{\Delta t,\theta}:=\tilde{Z}_{k}^{\theta}.$
Then, as $\Delta t\to 0$,
$\sup_{\theta\in\Theta_{0}}\sup_{t\in[0,T]}\mathbb{E}\Big[\,|p_{t}^{\Delta t,\theta}-p_{t}^{\theta}|^{2}+|p_{x,t}^{\Delta t,\theta}-p_{x,t}^{\theta}|^{2}+\|p_{y,t}^{\Delta t,\theta}-p_{y,t}^{\theta}\|^{2}\,\Big]$ $\to 0,$
$\sup_{\theta\in\Theta_{0}}\mathbb{E}\!\left[\int_{0}^{T}\Big(\|z_{t}^{\Delta t,\theta}-z_{t}^{\theta}\|^{2}+\|\tilde{z}_{t}^{\Delta t,\theta}-\tilde{z}_{t}^{\theta}\|^{2}\Big)\,dt\right]$ $\to 0.$
In particular, the convergence bounds can be taken uniform over $\theta$ in compact subsets of $\Theta$.
Proof sketch.
Fix $\theta\in\Theta$ and work conditionally on $\theta$ under the augmented simulator filtration $\mathcal{G}_{t}^{\theta}:=\sigma(\theta,\{W_{s},W_{s}^{Y}:0\leq s\leq t\})$ (usual augmentation). The argument follows the deterministic-parameter BPTT–BSDE/PMP correspondence of Huh et al. (2025), applied conditionally on $\theta$; uniformity on compact $\Theta_{0}\subset\Theta$ is obtained by assuming the required Lipschitz/growth/nondegeneracy/smoothness constants are uniform over $\theta\in\Theta_{0}$.
  1. 1.
    Pathwise BPTT adjoints on the Euler computation graph. For fixed $\theta$, simulate the Euler discretization of (39)–(40). BPTT computes the pathwise derivatives of the terminal utility along the full discrete computation graph, producing the pathwise adjoint blocks $(p_{k},p_{x,k},p_{y,k})$ as in (45)–(46). These quantities are generally $\mathcal{G}_{T}^{\theta}$-measurable (they depend on future increments) and hence are not adapted.
  2. 2.
    One-step conditional $L^{2}$ projection $\Rightarrow$ adapted discrete BSDE variables. On each step, project $p_{k+1}$ onto $\mathrm{span}\{1,\Delta W_{k},\Delta W_{k}^{Y}\}$ in conditional $L^{2}$ given $\mathcal{G}_{t_{k}}^{\theta}$:
    $p_{k+1}=\mathbb{E}[p_{k+1}\mid\mathcal{G}_{t_{k}}^{\theta}]+Z_{k}^{\theta}\,\Delta W_{k}+\tilde{Z}_{k}^{\theta}\,\Delta W_{k}^{Y}+\varepsilon_{k+1}^{\theta},$
    where $\varepsilon_{k+1}^{\theta}$ is orthogonal (in conditional $L^{2}$) to the span. Define the adapted discrete adjoints by conditional expectations,
    $p_{k}^{\Delta t,\theta}:=\mathbb{E}[p_{k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{x,k}^{\Delta t,\theta}:=\mathbb{E}[p_{x,k}\mid\mathcal{G}_{t_{k}}^{\theta}],\hskip 18.49988ptp_{y,k}^{\Delta t,\theta}:=\mathbb{E}[p_{y,k}\mid\mathcal{G}_{t_{k}}^{\theta}].$
    By construction, $(p_{k}^{\Delta t,\theta},Z_{k}^{\theta},\tilde{Z}_{k}^{\theta})$ satisfy a discrete-time BSDE representation.
  3. 3.
    Identification of the discrete driver; second-adjoint blocks and projection residual. Substituting the projection in Step 2 into the one-step BPTT recursion and taking conditional expectations yields that the discrete BSDE driver coincides with the Euler discretization of the $\theta$-conditional adjoint BSDE under the fixed policy $\pi_{\varphi}$, up to one-step remainder terms. In particular, in a smooth Markov regime (spelled out in Appendix D), the continuous-time adjoint admits a decoupling field $p_{t}^{\theta}=u^{\theta}(t,X_{t}^{\theta},Y_{t}^{\theta})$ and the sensitivity blocks needed in the portfolio Hamiltonian gradient coincide along trajectories with the spatial derivatives $\partial_{x}u^{\theta}$ and $\partial_{y}u^{\theta}$, yielding the interpretation of $(p_{x,t}^{\theta},p_{y,t}^{\theta})$. The projection residual $\varepsilon_{k+1}^{\theta}$ corresponds to higher-order chaos components and is controlled by the same one-step Taylor/chaos estimates as in Huh et al. (2025), contributing only higher-order discretization error (uniformly on compacts).
  4. 4.
    Convergence as $\Delta t\to 0$ and uniformity on compacts. Standard Euler stability for the forward SDE together with convergence of the associated discrete BSDE scheme yields, as $\Delta t\to 0$, mean-square convergence of the adapted discrete adjoint blocks $(p^{\Delta t,\theta},p_{x}^{\Delta t,\theta},p_{y}^{\Delta t,\theta})$ to $(p^{\theta},p_{x}^{\theta},p_{y}^{\theta})$ along the controlled trajectory, and convergence of the martingale coefficients $(Z^{\Delta t,\theta},\tilde{Z}^{\Delta t,\theta})$ to $(Z^{\theta},\tilde{Z}^{\theta})$ in the natural $L^{2}([0,T]\times\Omega)$ sense. If the coefficient/policy regularity and nondegeneracy constants are uniform for $\theta\in\Theta_{0}$, then the convergence bounds hold with constants independent of $\theta\in\Theta_{0}$.
A complete proof (including explicit regularity assumptions, the one-step remainder control, and the uniform-on-compacts bookkeeping) is provided in Appendix D. ∎
@@
Across $\theta\sim q$, these Pontryagin objects form a $\theta$-indexed family. Baseline DPO trains against the ex–ante objective (42) by repeatedly sampling $\theta\sim q$ inside the simulator, while the deployable policy remains $\theta$-blind.
@@
3.2 Stage 2 of PG–DPO under latent $\theta$: $q$-aggregated projection, diagnostics, and a residual-based policy-gap bound
@@
Stage 2 (Pontryagin guidance) is a projection step. Given a warm-up deployable $\theta$-blind feedback policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$ obtained from Stage 1 (DPO), we estimate $\theta$-conditional Pontryagin adjoint objects under frozen $\theta\sim q$ by BPTT/Monte Carlo, aggregate them across $\theta$, and construct a deployable $\theta$-blind policy by projecting onto the $q$-aggregated stationarity condition derived in Section 2.2. Together with Stage 1, this projection completes PG–DPO.
Equivalently, Stage 2 can be viewed as a plug-in approximation of one application of the population projection map $\mathcal{T}(\pi):=-A_{\pi}^{-1}g_{\pi}^{\mathrm{mix}}$ (defined below) evaluated at $\pi=\pi^{\mathrm{warm}}$ on the working domain. The key point is that the aggregated first-order condition is affine in the portfolio control; hence it induces a statewise linear system and, on a suitable working domain, a concrete projection map from (estimated) adjoint objects to a portfolio rule.
@@
Working domain and norms. Fix a measurable working state domain $\mathcal{D}\subset[0,T]\times(0,\infty)\times\mathbb{R}^{m}$ (e.g. a training/evaluation band) and a reference measure $\mu$ on $\mathcal{D}$ (e.g. an empirical state distribution induced by rollouts). For $h:\mathcal{D}\to\mathbb{R}^{n}$ we write
$\|h\|_{L^{2}(\mu)}:=\Big(\int_{\mathcal{D}}\|h(z)\|^{2}\,\mu(dz)\Big)^{1/2},\hskip 18.49988ptz=(t,x,y),$
@@
$\|f\|_{L^{2}(q\otimes\mu)}:=\bigg(\int_{\Theta}\int_{\mathcal{D}}\|f^{\theta}(z)\|^{2}\,\mu(dz)\,q(d\theta)\bigg)^{1/2}.$ (47)
@@
By Theorem 1, any locally optimal interior deployable $\theta$-blind policy $\pi^{\star,\mathrm{blind}}$ for the fixed-$q$ ex–ante problem satisfies the $q$-aggregated stationarity condition (18). In the portfolio Hamiltonian (11), this stationarity is equivalent to a statewise linear system and hence to the projected form (23) on the working domain (under invertibility of the aggregated curvature term). Stage 2 implements a practical approximation of this projection by estimating the relevant aggregated Pontryagin objects under the fixed warm-up policy $\pi^{\mathrm{warm}}=\pi_{\varphi^{\mathrm{warm}}}$.
Fix a query state $z=(t,x,y)\in\mathcal{D}$ and a frozen parameter $\theta$. We simulate trajectories under $\pi^{\mathrm{warm}}$ and compute discrete-time adjoint blocks by autodiff/BPTT on the Euler simulator; averaging over $M_{\mathrm{MC}}$ trajectories yields Monte Carlo estimates
@@
Whenever $\widehat{A}_{t}(t,x,y)$ is invertible on $\mathcal{D}$, we obtain the mixed-moment projected policy
@@
Stabilized inversion and projection diagnostics (implementation). In practice, finite-sample estimates can make $\widehat{A}_{t}$ ill-conditioned or even singular on parts of $\mathcal{D}$. Our default implementation therefore uses standard stabilizers (Appendix F.5), including: (i) ridge regularization $\widehat{A}_{t,\lambda}:=\widehat{A}_{t}+\lambda I_{d}$ and $\widehat{\pi}^{\mathrm{agg,mix}}_{\lambda}:=-\widehat{A}_{t,\lambda}^{-1}\widehat{g}_{t}^{\mathrm{mix}}$, (ii) condition-number or eigenvalue diagnostics to skip/clip projection on unreliable states (falling back to $\pi^{\mathrm{warm}}$), and (iii) residual-form projection (Section G.1) to reduce variance. The analysis below is stated for the population projection $\mathcal{T}(\pi):=-A_{\pi}^{-1}g_{\pi}^{\mathrm{mix}}$ and for small estimator perturbations, which is precisely the regime enforced by these stabilizers.
@@
To connect (53) to a locally optimal deployable $\theta$-blind policy, we measure how well the warm-up policy satisfies the population mixed-moment aggregated stationarity. Let $(A_{\pi},g_{\pi}^{\mathrm{mix}})$ denote the mixed-moment $q$-aggregated projection inputs induced by a policy $\pi$ (i.e. the objects in (21) evaluated using the $\theta$-conditional Pontryagin adjoint blocks generated by $\pi$). Define the warm-up aggregated stationarity residual on $\mathcal{D}$ by
$r_{\mathrm{FOC,mix}}^{\mathrm{warm}}(t,x,y):=A_{\pi^{\mathrm{warm}}}(t,x,y)\,\pi^{\mathrm{warm}}(t,x,y)+g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}(t,x,y),\hskip 18.49988pt\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}:=\big\|r_{\mathrm{FOC,mix}}^{\mathrm{warm}}\big\|_{L^{2}(\mu)}.$ (54)
Note that $r_{\mathrm{FOC,mix}}^{\mathrm{warm}}=0$ is equivalent to $\pi^{\mathrm{warm}}$ being a fixed point of $\mathcal{T}$ on $\mathcal{D}$. More generally, the population projection admits the residual form
$\mathcal{T}(\pi^{\mathrm{warm}})\;=\;\pi^{\mathrm{warm}}\;-\;A_{\pi^{\mathrm{warm}}}^{-1}\,r_{\mathrm{FOC,mix}}^{\mathrm{warm}}.$ (55)
In practice we monitor the empirical analogue $\widehat{r}_{\mathrm{FOC,mix}}^{\mathrm{warm}}:=\widehat{A}_{t}\,\pi^{\mathrm{warm}}+\widehat{g}_{t}^{\mathrm{mix}}$, and use it as a diagnostic to flag states where projection is unreliable (cf. Appendix F.5). (Equivalently, $\widehat{\pi}^{\mathrm{agg,mix}}=\pi^{\mathrm{warm}}-\widehat{A}_{t}^{-1}\widehat{r}_{\mathrm{FOC,mix}}^{\mathrm{warm}}$, or its stabilized variant, which motivates the residual-form implementation.)
@@
A technical point is that a global small-gain condition of the form $C_{1}<1$ can be overly restrictive. Following the slab-wise philosophy in our prior PGDPO analysis (e.g. Huh et al. (2025, Appendix B)), we default to a time-slab decomposition of the working domain and close the warm-up gap on each short slab. Concretely, assume $\mathcal{D}$ carries a time coordinate and fix a partition $0=t_{0}<t_{1}<\cdots<t_{L}=T$ with slab lengths $\tau_{k}:=t_{k}-t_{k-1}$. For notational convenience, write $\mathcal{D}\subset[0,T]\times\mathcal{S}$ for some spatial band $\mathcal{S}\subset(0,\infty)\times\mathbb{R}^{m}$, and define
$\mathcal{D}_{k}:=\mathcal{D}\cap([t_{k-1},t_{k}]\times\mathcal{S}),\hskip 18.49988pt\mu_{k}:=\mu|_{\mathcal{D}_{k}},\hskip 18.49988pt\|f\|_{k}:=\|f\|_{L^{2}(\mu_{k})}.$
We write $\mathcal{T}(\pi):=-A_{\pi}^{-1}g_{\pi}^{\mathrm{mix}}$ for the (population) $q$-aggregated projection map. Theorem 3 below shows that, under a mild slab-wise local stability regime (i.e. a short-time contraction of $\mathcal{T}$ on each $\mathcal{D}_{k}$), small residual implies that the projected policy is close (in $L^{2}(\mu)$) to a locally optimal deployable $\theta$-blind policy, up to discretization/Monte Carlo error. The proof combines a projection-map stability bound (Appendix E.1) with a slab-wise closure (Appendix E.2), in the same spirit as the slab analyses used in Huh et al. (2025).
Stage 1: DPO
@@
Stage 2: Pontryagin guidance
@@
Figure 1: Two-stage PG–DPO pipeline of Section 3: Stage 1 (DPO) trains a $\theta$-blind policy by BPTT, and Stage 2 applies a $q$-aggregated Pontryagin projection. A detailed schematic (including coupling mechanisms and implementation notes) is provided in Appendix G, Fig. 4.
Theorem 3 (Residual-based ex–ante $\theta$-blind policy-gap bound for PG–DPO (mixed-moment, deployable, slab-wise local)).
Let $\pi^{\mathrm{warm}}$ be the Stage 1 (DPO) warm-up policy and let $(\widehat{A}_{t},\widehat{g}_{t}^{\mathrm{mix}})$ be the Stage 2 mixed-moment estimators computed under $\pi^{\mathrm{warm}}$. Assume Proposition 1 (Appendix E.1) applies on the working domain $\mathcal{D}$ with $A=A_{\pi^{\mathrm{warm}}}$ and $g=g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}$; equivalently, there exist constants $\kappa,M_{g}>0$ such that
$\|A_{\pi^{\mathrm{warm}}}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa,\hskip 18.49988pt\|g_{\pi^{\mathrm{warm}}}^{\mathrm{mix}}\|_{L^{\infty}(\mathcal{D})}\leq M_{g},\hskip 18.49988pt\|\widehat{A}_{t}-A_{\pi^{\mathrm{warm}}}\|_{L^{\infty}(\mathcal{D})}\leq(2\kappa)^{-1},$
where the last condition is the stable inversion regime enforced in practice by ridge/diagnostics/skip.
Let $\pi^{\star,\mathrm{blind}}$ be a locally optimal interior deployable $\theta$-blind policy for the fixed-$q$ ex–ante problem. Assume there exists a neighborhood $\mathcal{U}$ of $\pi^{\star,\mathrm{blind}}$ in $L^{2}(\mu)$ such that:
  1. (B1)
    $\pi^{\mathrm{warm}}\in\mathcal{U}$ and, for all $\pi\in\mathcal{U}$, $\|A_{\pi}^{-1}\|_{L^{\infty}(\mathcal{D})}\leq\kappa$ and $\|g_{\pi}^{\mathrm{mix}}\|_{L^{\infty}(\mathcal{D})}\leq M_{g};$
  2. (B2)
    the slab-wise Lipschitz gain of Appendix E.2 holds: there exist constants $\bar{L}_{A},\bar{L}_{g}>0$ such that for every slab $\mathcal{D}_{k}$ and all $\pi_{1},\pi_{2}\in\mathcal{U}$,
@@
$\varrho(\tau):=\big(\kappa\bar{L}_{g}+\kappa^{2}M_{g}\bar{L}_{A}\big)\tau^{1/2},\hskip 18.49988pt\varrho_{*}:=\max_{1\leq k\leq L}\varrho(\tau_{k}),$ (58)
and assume the slab partition is chosen so that $\varrho_{*}<1$.
@@
$\big\|\widehat{\pi}^{\mathrm{agg,mix}}-\pi^{\star,\mathrm{blind}}\big\|_{L^{2}(\mu)}\;\leq\;\frac{\varrho_{*}\kappa}{1-\varrho_{*}}\,\varepsilon_{\mathrm{warm}}^{\mathrm{mix}}\;+\;C_{2}\,\delta_{\mathrm{BPTT}}(\Delta t,M_{\mathrm{MC}},M_{\theta}),$ (59)